Westerville Company reported the following results from last year’s operations:

Sales- 1,600,000

Variable expenses-700,000

Contribution margin-900,000

Fixed expenses-660,000

Net operating income-240,000

Average operating assets-1,000,000

At the beginning of this year, the company has a $325,000 investment opportunity with the following cost and revenue characteristics:

Sales-520,000

Contribution margin ratio-70% of sales

Fixed expenses-312,000

The company’s minimum required rate of return is 15%.

A) What is last year’s margin?

B) What is last year’s return on investment (ROI)?

C) What is the margin related to this year’s investment opportunity?

D)What is the ROI related to this year’s investment opportunity?

E) If the company pursues the investment opportunity and otherwise performs the same as last year, what margin will it earn this year?

F) If the company pursues the investment opportunity and otherwise performs the same as last year, what turnover will it earn this year?

G) If the company pursues the investment opportunity and otherwise performs the same as last year, what ROI will it earn this year?

H) What is last year’s residual income?

I) What is the residual income of this year’s investment opportunity?

J) . If the company pursues the investment opportunity and otherwise performs the same as last year, what residual income will it earn this year?
Answer #1

dear student,

as per the HomeworkLib policy, only the first four questions should be answered. kindly take note of it.

Part A

last year’s margin = Net operating income/ sales = 240000/1600000 = 15%

Part B

last year’s return on investment (ROI) = margin * turnover

turnover = sales / average operating assets = 1600000/1,000,000 = 1.6

last year’s return on investment (ROI) = 15%*1.6 = 24%

Part C

margin for this year’s investment opportunity = Net operating income/ sales = ((520000*70%)-312000)/520000 = 10%

Part D

ROI related to this year’s investment opportunity = margin * turnover

turnover = sales / average operating assets = 520000/325000 = 1.60

ROI related to this year’s investment opportunity = 10%*1.6 = 16%